# Fair Uncertainty Quantification for Depression Prediction

### v2

@@ Abstract
Trustworthy depression prediction based on deep learning, incorporating both predictive reliability and algorithmic fairness across diverse demographic groups, is crucial for clinical application. Recently, achieving reliable depression predictions through uncertainty quantification has attracted increasing attention. However, few studies have focused on the fairness of uncertainty quantification (UQ) in depression prediction. In this work, we investigate the algorithmic fairness of UQ, namely Equal Opportunity Coverage (EOC) fairness, and propose Fair Uncertainty Quantification (FUQ) for depression prediction. FUQ pursues reliable and fair depression predictions through group-based analysis. Specifically, we first group all the participants by different sensitive attributes and leverage conformal prediction to quantify uncertainty within each demographic group, which provides a theoretically guaranteed and valid way to quantify uncertainty for depression prediction and facilitates the investigation of fairness across different demographic groups. Furthermore, we propose a fairness-aware optimization strategy that formulates fairness as a constrained optimization problem under EOC constraints. This enables the model to preserve predictive reliability while adapting to the heterogeneous uncertainty levels across demographic groups, thereby achieving optimal fairness. Through extensive evaluations on several visual and audio depression datasets, our approach demonstrates its effectiveness.
Index Terms:
Depression prediction, uncertainty quantification, algorithmic fairness, conformal prediction.
@@ I Introduction
Figure 1: Comparison of uncertainty quantification (UQ) approaches for depression prediction: (a) Existing UQ provides statistically valid confidence intervals globally, but exhibits disparate coverage rates across demographic groups (highlighted in bold) [11]. (b) Fair Uncertainty Quantification (FUQ) achieves dual objectives: maintaining global statistical validity (comparable to existing UQ) while ensuring equitable coverage across all subgroups. The bold-highlighted group-wise coverage rates demonstrate FUQ’s fairness improvement.
@@
In this context, deep learning-based Automatic Depression Recognition Systems (ADRS) have emerged as a promising solution by integrating multimodal behavioral data, including speech prosody, facial expression dynamics, and motor biomarkers, to enable a quantitative analysis of depressive traits [16, 17, 18, 19, 20, 21]. These systems capture depression-related features in naturalistic interaction settings, overcoming the spatial and temporal constraints of traditional diagnostic methods. Moreover, ADRS facilitate the development of dynamic monitoring frameworks, providing a transformative paradigm for early detection and timely intervention in depression.
@@
As research delves deeper, despite the outstanding capabilities for learning representations, researchers have found that the deep learning models are untrustworthy. One major reason is that deep models are typically ”black boxes” and lack interpretability. Additionally, observations have shown that predictions often exhibit ”overconfidence”. In depression prediction, [22, 11] address this issue through uncertainty quantification (UQ). Calibrated predictions with UQ not only provide accurate predictions but also output predictions with the confidence level.
@@
Moreover, few studies have focused on the issue that data-driven deep learning algorithms are always prone to predictive bias, which varies across different groups and leads to unfair predictions. Such biases can come from various sources, including imbalanced data distribution, inherent algorithmic design, and task-specific requirements [23, 24, 25]. Kuzucu et al. [26] show that models produce systematically biased uncertainty estimates across demographic groups. Furthermore, empirical evidence shows that depression prediction is unfair across demographic groups, and such bias directly weakens the validity of uncertainty quantification, as shown in Fig. 1(a). [27, 28] seeks to achieve fair depression prediction by eliminating biases associated with specific attributes, but they overlook uncertainty quantification and calibration. However, achieving fair depression prediction within a UQ framework is nontrivial [26]. First, incorporating fairness considerations into existing UQ methods for depression recognition is neither intuitive nor straightforward. Most prevailing approaches prioritize aggregate statistical validity and, in doing so, neglect the demographic stratifications with sensitive attributes. Consequently, their uncertainty estimates tend to privilege majority groups while failing to adequately cover minority groups, thereby introducing unfairness. Second, stratifying the data by sensitive attributes substantially reduces the sample size within each group, which can, in turn, impair the statistical validity of the resulting UQ estimates. In this paper, we consider reliable depression prediction through effective UQ and fairness-aware optimization. The motivation for this work is summarized as follows:
@@
    Uncertainty quantification in depression prediction is inherently influenced by algorithmic design and calibration.
@@
Given the absence of a well-established fairness definition in depression prediction, we focus on Equal Opportunity Coverage (EOC) fairness, an uncertainty-based fairness notion introduced in [29]. As illustrated in Fig. 1, compared to previous uncertainty quantification methods, EOC requires that the validity of UQ should not only satisfy the desired confidence rate across all participants but also maintain comparable validity across demographic groups (e.g., male and female). Our contributions can be summarized as follows:
1. We propose Fair Uncertainty Quantification (FUQ), a framework for achieving fair and reliable predictions through group-based analysis. To the best of our knowledge, this work is the first to explore EOC-fairness in depression prediction through the lens of uncertainty quantification.
2. In FUQ, we design a fairness-aware optimization strategy to preserve EOC fairness across demographic groups, while providing theoretically guaranteed and calibrated uncertainty estimates for each group.
@@ II-A Depression Recognition
Depression recognition has attracted significant attention, initially originating from the AVEC challenge, which used the AVEC 2013 and 2014 facial video datasets to inspire a considerable community of researchers [30, 31]. Early efforts, such as those by Williamson et al., employed handcrafted audio features and trained a multivariate regression Gaussian mixture model (GMM), achieving notable success in the competition [16].
In recent years, deep learning has largely replaced traditional handcrafted features, leveraging its powerful representational capabilities to enhance recognition accuracy. Melo et al. further advanced the field by proposing the Multiscale Spatiotemporal Network (MSN) and Maximization-Differentiation Network (MDN) to improve feature extraction for a more comprehensive depression representation [21, 20]. Zhang et al. designed the lightweight Multi-scale Temporal Difference Attention Network (MTDAN) to address the high computational complexity of deep learning models, facilitating real-time clinical applications [32]. Niu et al. estimated depression severity using Facial Keypoints Representation Sequences (FKRS) and Action Units Representation Sequences (AURS), avoiding the use of raw facial video data [33]. Xu et al. proposed automatic facial depression recognition based on a two-stage framework, where short-term facial behaviors are first modeled, followed by video-level depression analysis using Sequential Graph Representation (SEG) and Spectral Graph Representation (SPG) [34].
In addition to using a single modality, some studies have also focused on multimodal depression recognition. Lang et al. applied deep neural networks (DNNs) for multimodal fusion of audio, video, and text [35]. Mao et al. developed an attention-based multimodal representation by integrating Bi-LSTM with GloVe embeddings [18], while Uddin et al. introduced a video-speech multimodal prediction model that employs Volume Local Directional Structural Pattern (VLDSP-based) dynamic feature descriptors for facial dynamics and Temporal Attention Pooling (TAP) for segment-level audiovisual feature aggregation [36]. Pan et al. proposed an audiovisual attention network (AVA-DepressNet), a multimodal framework that enhances audiovisual spatial-temporal features through attention mechanisms and optimizes the encoder-decoder structure using an adversarial multistage training strategy with facial structural priors [17]. Fan et al. further extended multimodal depression detection by introducing Transformer-based Multimodal Feature Enhancement Networks (TMFE) and employing a graph fusion network architecture to integrate features across different modalities [37]. While advancements have reduced errors in depression recognition, the above methods overlook uncertainty and predictive bias, thereby undermining the trustworthiness of their predictions. Addressing these challenges remains crucial for developing trustworthy ADRS.
@@ II-B Uncertainty and Bias in Affective Computing
In the field of affective computing, UQ has attracted significant attention. Zhou et al. utilized label distributions to learn the uncertainty of predictions, where the smoother the learned label distribution, the greater the uncertainty [38]. Lo et al. modeled uncertainty in low-resolution facial expressions through structured probabilistic embeddings [39]. Ahn et al. enhanced representation learning by combining contrastive learning with features represented as probability distributions to model uncertainty [40]. Li et al. modeled uncertainty in depression prediction through interval predictions, where wider intervals indicate higher uncertainty [11]. These methods incorporate uncertainty analysis into affective computing through heuristic approaches or uncertainty quantification. However, these methods prioritize the statistical performance of UQ and overlook predictive biases across demographic groups, resulting in inconsistent quantification performance.
@@
On the other hand, data-driven affective computing methods are often prone to data or algorithmic bias. Gratch et al. highlighted ethical considerations in affective computing, emphasizing the biases arising from insufficient diversity in dataset attributes [41, 42]. Similar observations are evident in related studies. Li et al. investigated biases in facial expression recognition datasets and proposed the Emotion-Conditional Adaptation Network (ECAN) to learn domain-invariant facial representations [43]. Adarsh et al. combined Support Vector Machines (SVM) and K-Nearest Neighbors (KNN) in an ensemble model to address age-related prediction biases, striving for fair depression detection from social media data [44]. Additionally, Dang et al. addressed prediction unfair biases in depression prediction using machine learning, proposing Population Sensitivity-Guided Threshold Adjustment (PSTA) to achieve fair prediction [45]. Ensuring fairness is essential for developing ADRS that are not only accurate but also trustworthy and ethically responsible. Notably, Cheong et al. introduced the FAIRREFUSE framework, a dynamic referee-guided causal approach employing causal intervention and individual fairness-guided fusion to mitigate gender biases in multimodal depression recognition [27]. Additionally, they proposed a novel gender-based task-reweighting approach leveraging uncertainty estimation to enhance performance and fairness in depression detection [28].
@@
Prior studies have addressed predictive bias in depression recognition and have maintained fairness by mitigating such bias. Nevertheless, incorporating fairness into UQ remains challenging. Traditional UQ performance metrics emphasize overall statistical validity, which depends on aggregate sample size and therefore tend to overlook disparities associated with sensitive attributes; furthermore, stratifying data by these attributes reduces group sample sizes and consequently degrades UQ validity. Cheong et al. share a similar underlying motivation with ours [28], concentrating on fairness-aware depression detection via uncertainty-driven heuristics formulated as a classification task. By contrast, our study integrates fairness into rigorously calibrated UQ and advances reliable depression prediction formulated as a regression task.
Figure 2: The pipeline of FUQ. FUQ achieves the validity of prediction intervals through group conditional conformal prediction, while allowing independent optimization across different group. By further employing a fairness-aware optimization strategy, FUQ guarantees that its uncertainty quantification is both statistically valid and fair across demographic groups.
@@
$\hat{q}$: well-trained quantile model on the training set;$\mathcal{Q}(\cdot,\cdot)$: quantile estimator;
@@
$B_{1},B_{2},...,B_{M}$: $M$ equal-mass bins; $l_{m},u_{m}$: lower and upper bounds of $B_{m}$; $|B_{m}|$: sample sizes in the bin $B_{m}$.
0:  Fair and reliable depression prediction $C(X_{test}^{s})$.
1:  Initialize sub-coverage $\beta_{m,s}=0$// Calculate the $1-\alpha$ intervals of CQR
2:  for $i=1$ to $N$ do
3:     $r_{i}=\max\{\hat{q}_{\alpha/2}(X_{i}^{s})-y_{i},y_{i}-\hat{q}_{1-\alpha/2}(X_{i}^{s})\}$
4:  end for
5:  $\hat{r}=\mathcal{Q}(1-\alpha,r)$
6:  for $i=1$ to $N$ do
7:     $C(X_{i}^{s})=[\hat{q}_{\alpha/2}(X_{i}^{s})-\hat{r},\hat{q}_{1-\alpha/2}(X_{i}^{s})+\hat{r}]$ // Calculate the coverage rate for each sensitive attributes within each bin
8:     for $m=1$ to $M$ do
9:        for $s^{\prime}=1$ to $S$ do
10:           $\beta_{m,s}\leftarrow\beta_{m,s}+\frac{\mathbf{1}\{y_{i}\in(B_{m}\cap C(X_{i}^{s}))\}*\mathbf{1}\{s=s^{\prime}\}}{|B_{m}|}$
@@
13:  end for// Using Algorithm 2 to obtain $\hat{r}_{m,s}$
@@
15:     $C_{m}(X_{test}^{s})=[\hat{q}_{\alpha/2}(X_{test}^{s})-\hat{r}_{m,s},\hat{q}_{1-\alpha/2}(X_{test}^{s})+\hat{r}_{m,s}]\cap[l_{m},u_{m})$
@@
III Method
III-A Preliminaries
Let the depression dataset be divided into a training set $\mathcal{D}_{train}$, a calibration set $\mathcal{D}_{cal}$, and a test set $\mathcal{D}_{test}$. We denote $s$ as the sensitive attribute of the input $X$ (such as age, gender), $S$ as the total number of attributes, and $y$ as the depression score. On the $\mathcal{D}_{train}$, we can achieve the vanilla depression prediction $\hat{y}=f(X^{s})$ by minimizing the MSE loss as the learning objective of the deep neural network $f(\cdot)$.
III-B UQ for depression prediction
Referring to [11], we adopt Conformalized Quantile Regression (CQR) [46] for uncertainty quantification in depression prediction. Specifically, during training, we apply the pinball loss function as the learning objective to achieve quantile prediction for depression scores.
where $q\in[0,1]$, $\hat{y}_{q}$ denotes the $q$-th quantile of the prediction. This enables the model to learn the upper and lower quantile bounds, allowing for the construction of prediction intervals that represent uncertainty effectively. Let $\alpha$ denote the miscoverage rate. On the $\mathcal{D}_{cal}$, we then utilize the ($1-\alpha$)-th $\max\{\hat{y}_{\alpha/2}-y,y-\hat{y}_{1-\alpha/2}\}$} to obtain $\beta$, the $1-\alpha$ quantile of the conformal scores. The valid $1-\alpha$ confidence interval is denoted as $C(X)=[\hat{y}_{\alpha/2}-\beta,\hat{y}_{1-\alpha/2}+\beta]$ for $\mathcal{D}_{test}$.
III-C FUQ for depression prediction
Through CQR, we acquire calibrated confidence intervals that statistically ensure $1-\alpha$ coverage for all the participants. However, empirical observations reveal that coverage varies across sensitive attribute groups, indicating non-uniform effectiveness. Some groups demonstrate over-coverage, while others do not meet the required coverage guarantee. Therefore, we pursue Equal Opportunity Coverage (EOC):
$P(y\in C(X^{s})|s=s_{0})=P(y\in C(X^{s})|s=s_{1})\geq 1-\alpha$ (2)
Achieving Equal Opportunity Coverage (EOC) entails two conditions. First, the coverage across different sensitive attribute groups should be equal (we illustrate with groups $s_{0}$ and $s_{1}$, though practical applications may involve more). Second, each group should meet the expected coverage. The most straightforward method is to group the data based on sensitive attributes and apply CQR to output confidence intervals separately within each group. While this sounds theoretically promising, grouping significantly sharpens the available data for each group, thereby diminishing the statistical validity guarantees of UQ.
We leverage a group conditional conformal prediction with equal-mass binning to maximize within-group validity and to facilitate the implementation of fairness interventions [29]. Specifically, it divides the continuous variable $y$ into $M$ bins with equal sample sizes, $B_{m}=[l_{m},u_{m})$, $m\in[1,M]$, and $l,u$ denote the lower and upper bounds of the bin, respectively. This bin-dividing method ensures that each bin is given equal importance while maximizing the size of each bin, and it is beneficial to permit the independent optimization of coverage for $M$ bins, thereby ensuring marginal coverage while maintaining the interval widths at their narrowest possible point. Since each bin in group conditional conformal prediction is optimized independently, we still need a fairness-aware optimization for the prediction interval while keeping the interval widths as narrow as possible to be informative. That is
$\min$ $\sum_{m,i\in|\mathcal{D}_{t}|}\frac{|C_{m}(X_{i})|}{|\mathcal{D}_{cal}|}$ (3)
s.t. $\begin{cases}\sum_{m}\frac{\beta_{m,s}}{M}=1-\alpha,\quad s\in[1,S],\\ \beta_{m,s}\in[0,1],\quad m=1,\ldots,M.\end{cases}$
By leveraging the gradient of coverage with respect to conformal scores, we optimize the confidence intervals across different attributes, resulting in fairer UQ for depression prediction, shown in Algorithm 1.

### v3

@@ Abstract
Trustworthy depression prediction based on deep learning, incorporating both predictive reliability and algorithmic fairness across diverse demographic groups, is crucial for clinical applications. Recently, achieving reliable depression predictions through uncertainty quantification has attracted increasing attention. However, few studies have focused on the fairness of uncertainty quantification (UQ) in depression prediction. In this work, we investigate the algorithmic fairness of UQ, namely equal opportunity of coverage (EOC) fairness, and propose fair uncertainty quantification (FUQ) for depression prediction. FUQ pursues reliable and fair depression predictions through group-based analysis. Specifically, we first group all the participants by different sensitive attributes and leverage the group conditional conformal prediction method to quantify uncertainty within each demographic group, which provides a theoretically guaranteed and valid way to quantify uncertainty for depression prediction and facilitates the investigation of fairness across different demographic groups. Furthermore, we formulate the fairness of UQ as a constrained optimization problem under EOC constraints, and propose a fairness-aware optimization strategy to solve it. This enables the model to preserve predictive reliability while adapting to the heterogeneous uncertainty levels across demographic groups, thereby achieving optimal fairness. Through extensive evaluations on several visual and audio depression datasets, our approach demonstrates its effectiveness.
@@ I Introduction
Figure 1: Comparison of uncertainty quantification (UQ) approaches for depression prediction (using gender as the sensitive attribute): (a) Existing UQ provides statistically valid confidence intervals globally for All participants, but exhibits disparate coverage rates across demographic groups for gender. (b) Fair uncertainty quantification (FUQ) achieves dual objectives: maintaining global statistical validity while ensuring equitable coverage across all subgroups. Here, $B$ indicates bins formed by evenly dividing participants according to depression level; this binning supports approximate conditional coverage in our interval construction.
@@
In this context, deep learning-based Automatic Depression Recognition systems (ADRs) have emerged as a promising solution by behavioral data, including speech prosody, facial expression dynamics, and motor biomarkers, to enable a quantitative analysis of depressive traits [16, 17, 18, 19, 20, 21]. These systems capture depression-related features in naturalistic interaction settings, overcoming the spatial and temporal constraints of traditional diagnostic methods. Moreover, ADRs facilitate the development of dynamic monitoring frameworks, providing a transformative paradigm for early detection and timely intervention in depression.
@@
As research delves deeper, despite the outstanding capabilities for learning representations, researchers have found that the deep learning models are often untrustworthy [22]. One major reason is that deep models are typically black boxes and lack interpretability. Additionally, observations have shown that predictions often exhibit overconfidence. In depression prediction, [23, 11] address this issue through uncertainty quantification (UQ). Calibrated predictions with UQ not only provide accurate predictions but also output predictions with the confidence level.
@@
Moreover, few studies have focused on the issue that data-driven deep learning algorithms are always prone to predictive bias, which varies across different groups and leads to unfair predictions. Such biases can come from various sources, including imbalanced data distribution, inherent algorithmic design, and task-specific requirements [24, 25, 26]. Kuzucu et al. [27] show that models produce systematically biased uncertainty estimates across demographic groups. Furthermore, empirical evidence shows that depression prediction is unfair across demographic groups, and such bias directly weakens the validity of uncertainty quantification, as shown in Fig. 1(a). [28, 29] seeks to achieve fair depression prediction by eliminating biases associated with specific attributes, but they overlook uncertainty quantification and calibration. Ensuring fairness in UQ is as critical as UQ itself; unfair uncertainty estimates can impose disproportionate and potentially catastrophic harms on a specific group. For example, if uncertainty is well calibrated for younger patients with depression but neglected for older patients, it would undermine trustworthiness for the latter and could lead to severe, irreversible harm.
However, achieving fair depression prediction within a UQ framework is nontrivial [27]. First, incorporating fairness considerations into existing UQ methods for depression recognition is neither intuitive nor straightforward. Most prevailing approaches prioritize aggregate statistical validity and, in doing so, neglect the demographic stratifications with sensitive attributes. Consequently, their uncertainty estimates tend to privilege majority groups while failing to adequately cover minority groups, thereby introducing unfairness. Second, stratifying the data by sensitive attributes substantially reduces the sample size within each group, which can, in turn, impair the statistical validity of the resulting UQ estimates. In this paper, we consider reliable depression prediction through effective UQ and fairness-aware optimization. The motivation for this work is summarized as follows:
@@
    Uncertainty quantification in depression prediction is inherently influenced by algorithmic design and uncertainty calibration.
@@
Given the absence of a well-established fairness definition in depression prediction, we focus on equal opportunity of coverage (EOC) fairness, an uncertainty-based fairness notion introduced in [30]. As illustrated in Fig. 1, compared to previous uncertainty quantification methods, EOC requires that the validity of UQ should not only satisfy the desired confidence rate across all participants but also maintain comparable validity across demographic groups (e.g., male and female). Our contributions can be summarized as follows:
1. We propose fair uncertainty quantification (FUQ), a framework for achieving fair and reliable depression predictions through group-based analysis. To the best of our knowledge, this work is the first to explore EOC-fairness in depression prediction through the lens of uncertainty quantification.
2. In FUQ, we design a fairness-aware optimization strategy to preserve EOC fairness in depression prediction across demographic groups, while providing theoretically guaranteed and calibrated uncertainty estimates for each group.
@@
Figure 2: The pipeline of FUQ for depression prediction. FUQ produces prediction interval that is both statistically valid and fair across demographic groups. It calibrates traditional prediction intervals (produced by quantile regression, for example) into fair intervals by coupling group conditional conformal prediction (ensuring reliable uncertainty estimates within group), with a fairness-aware optimization step (balancing inter-group coverage). $f$ denotes the quantile regression model, which outputs prediction intervals for uncertainty quantification (UQ), $M$ denotes the number of the dividing bins for conditional CP, and $S$ denotes the number of the sensitive attributes for grouping.
@@ II-A Depression Recognition
Depression recognition has attracted significant attention, initially originating from the AVEC challenge, which used the AVEC 2013 and 2014 facial videos to inspire a considerable community of researchers [31, 32]. Early efforts, such as those by Williamson et al., employed handcrafted audio features and trained a multivariate regression Gaussian mixture model (GMM), achieving notable success in the competition [16].
In recent years, deep learning has largely replaced traditional handcrafted features, leveraging its powerful representational capabilities to enhance recognition accuracy. Melo et al. advanced the field by proposing the multiscale spatiotemporal network (MSN) and maximization-differentiation network (MDN) to improve feature extraction for a more comprehensive depression representation [21, 20]. Zhang et al. designed the lightweight multiscale temporal difference attention network (MTDAN) to address the high computational complexity of deep learning models, facilitating real-time clinical applications [33]. Niu et al. estimated depression severity using facial keypoints representation sequences (FKRS) and action units representation sequences (AURS), avoiding the use of raw facial video data [34]. Xu et al. proposed automatic facial depression recognition based on a two-stage framework, where short-term facial behaviors are first modeled, followed by video-level depression analysis using sequential graph representation (SEG) and spectral graph representation (SPG) [35].
In addition to using a single modality, some studies have also focused on multimodal depression recognition. Yang et al. applied deep neural networks (DNNs) for multimodal fusion of audio, video, and text [36]. Mao et al. developed an attention-based multimodal representation by integrating Bi-LSTM with GloVe embeddings [18], while Uddin et al. introduced a video-speech multimodal prediction model that employs volume local directional structural pattern (VLDSP-based) dynamic feature descriptors for facial dynamics and temporal attention pooling (TAP) for segment-level audiovisual feature aggregation [37]. Pan et al. proposed an audiovisual attention network (AVA-DepressNet), a multimodal framework that enhances audiovisual spatial-temporal features through attention mechanisms and optimizes the encoder-decoder structure using an adversarial multistage training strategy with facial structural priors [17]. Fan et al. further extended multimodal depression detection by introducing transformer-based multimodal feature enhancement networks (TMFE) and employing a graph fusion network architecture to integrate features across different modalities [38]. While advancements have reduced errors in depression recognition, most existing methods overlook predictive uncertainty and predictive bias, thereby undermining the trustworthiness of their predictions. Addressing these challenges remains crucial for developing trustworthy ADRs.
@@ II-B Uncertainty and Bias in Affective Computing
In the field of affective computing, UQ has also attracted significant attention. Zhou et al. utilized label distributions to model the uncertainty of predictions, where the smoother the learned label distribution, the greater the uncertainty [39]. Lo et al. modeled uncertainty in low-resolution facial expressions through structured probabilistic embeddings [40]. Ahn et al. enhanced representation learning by combining contrastive learning with features represented as probability distributions to model uncertainty [41]. Li et al. modeled uncertainty in depression prediction through interval predictions, where wider intervals indicate higher uncertainty [11]. These methods incorporate uncertainty analysis into affective computing through either heuristic approaches or uncertainty quantification. However, these methods prioritize the statistical performance of UQ and overlook predictive biases across demographic groups, resulting in inconsistent quantification performance.
@@
On the other hand, data-driven affective computing methods are often prone to data or algorithmic bias. Gratch et al. highlighted ethical considerations in affective computing, emphasizing the biases arising from insufficient diversity in dataset attributes [42, 43]. Similar observations are evident in related studies. Li et al. investigated biases in facial expression recognition and proposed the emotion-conditional adaptation network (ECAN) to learn domain-invariant facial representations [44]. Adarsh et al. combined support vector machines (SVM) and K-Nearest Neighbors (KNN) in an ensemble model to address age-related prediction biases, striving for fair depression detection from social media data [45]. Additionally, Dang et al. addressed prediction unfair biases in depression prediction using machine learning, proposing population sensitivity-guided threshold adjustment (PSTA) to achieve fair prediction [46]. Ensuring fairness is essential for developing ADRs that are not only accurate but also trustworthy and ethically responsible. Notably, Cheong et al. introduced the FAIRREFUSE framework, a dynamic referee-guided causal approach employing causal intervention and individual fairness-guided fusion to mitigate gender biases in multimodal depression recognition [28]. Additionally, they proposed a gender-based task-reweighting approach leveraging uncertainty estimation to enhance performance and fairness in depression detection [29].
@@
Prior studies have addressed predictive bias in depression recognition and have maintained fairness by mitigating such bias. Nevertheless, incorporating fairness into UQ remains challenging. Traditional UQ performance metrics emphasize overall statistical validity, which depends on aggregate sample size and therefore tend to overlook disparities associated with sensitive attributes; furthermore, stratifying data by these attributes reduces group sample sizes and consequently degrades UQ validity. Cheong et al. share a partially similar underlying motivation with ours [29], concentrating on fairness-aware depression detection via uncertainty-driven heuristics formulated as a classification task. By contrast, our study integrates fairness into rigorously calibrated UQ, and advances reliable depression prediction formulated as a regression task.
III Method
III-A UQ for Depression Prediction
Let the depression dataset be divided into a training set $\mathcal{D}_{train}$, a calibration set $\mathcal{D}_{cal}$, and a test set $\mathcal{D}_{test}$. We denote $s$ as the sensitive attribute (such as age, gender) of the input $X$, and $y$ as the depression score. Note that $S$ denotes the number of categories for a given attribute (e.g., for gender, $S=2:$ male and female), not the number of distinct attributes. For the traditional depression prediction, we train a model by minimizing the loss (such as MAE, MSE) as the learning objective on the $\mathcal{D}_{train}$. However, the model does not incorporate uncertainty quantification capabilities. This work trains a quantile regression model to perform preliminary uncertainty quantification. Specifically, during training, we apply the pinball loss as the learning objective to achieve quantile prediction.
where $q\in[0,1]$, $\hat{y}_{q}$ denotes the $q$-th quantile of the prediction. This enables the model to learn the lower and upper quantile bounds $\hat{y}_{l}$ and $\hat{y}_{h}$, allowing for the construction of prediction intervals that represent uncertainty. To further calibrate the performance of uncertainty quantification, and following [11], we adopt conformalized quantile regression (CQR) [47] for uncertainty quantification in depression prediction. Let $\alpha$ denote the miscoverage rate. On the $\mathcal{D}_{cal}$, we then utilize the ($1-\alpha$)-th $\max\{\hat{y}_{\alpha/2}-y,y-\hat{y}_{1-\alpha/2}\}$} to obtain $\beta$, the $1-\alpha$ quantile of the conformal scores. The valid $1-\alpha$ confidence interval is denoted as $C(X)=[\hat{y}_{\alpha/2}-\beta,\hat{y}_{1-\alpha/2}+\beta]$ for $X\in\mathcal{D}_{test}$. CQR yields a ($1-\alpha$) level confidence interval $C$ with a marginal coverage guarantee, under exchangeability,
$P(y\in C(X))\geq 1-\alpha$ (2)
While valid, marginal coverage can be further refined to better approximate conditional coverage (even though [48] shows that conditional coverage is hard to achieve, it can be approximated).
$P(y\in C(X)|X=X_{i})\geq 1-\alpha$ (3)
The current limitation is that CQR uses the fixed global calibration factor $\beta$ that does not vary with a test sample’s depression severity. To approximate conditional coverage, we adopt a simple scheme by partitioning participants into $M$ equal-mass bins by depression severity, $B_{m}=[l_{m},u_{m})$, $m\in[1,M]$, and $l,u$ denote the lower and upper bounds of the bin respectively, and computing a group-specific calibration factor $\beta_{m}$ for each group. Thus, the prediction intervals can be regarded as achieving approximate conditional coverage across different levels of depression severity. However, empirical observations reveal that coverage varies across sensitive attribute groups, indicating non-uniform effectiveness. Some groups demonstrate over-coverage, while others do not meet the required coverage guarantee. Therefore, we pursue equal opportunity of coverage (EOC).
@@ III-A UQ for Depression Prediction
[…]
@@
$B_{1},B_{2},...,B_{M}$: $M$ equal-mass bins; $l_{m},u_{m}$: lower and upper bounds of $B_{m}$; $X_{test}^{s}$: test data with a specific attribute $s$.
0: Fair and reliable prediction $C(X_{test}^{s})$.// Calculate the $1-\alpha$ intervals of CQR and initialize subgroup coverage $\beta_{m,s}$
1:for $i=1$ to NN do
2:  $r_{i}=\max\{f_{\alpha/2}(X_{i}^{s})-y_{i},y_{i}-f_{1-\alpha/2}(X_{i}^{s})\}$
3:end for
[…]
5:for $i=1$ to NN do
6:  $C(X_{i}^{s})=[f_{\alpha/2}(X_{i}^{s})-\hat{r},f_{1-\alpha/2}(X_{i}^{s})+\hat{r}$]
7:  for $m=1$ to $M$ do
8:   for $s^{\prime}=1$ to $S$ do
9:    $\beta_{m,s}\leftarrow\frac{\mathbf{1}\{y_{i}\in(B_{m}\cap C(X_{i}^{s}))\}*\mathbf{1}\{s=s^{\prime}\}}{|B_{m}|}$
10:   end for
@@
13: Using Eqs. 6-11 to obtain $\hat{r}_{m,s}$ // Calculate the confidence interval
@@
15:  $C_{m}(X_{test}^{s})=[f_{\alpha/2}(X_{test}^{s})-\hat{r}_{m,s},f_{1-\alpha/2}(X_{test}^{s})+\hat{r}_{m,s}]\cap[l_{m},u_{m})$
@@
III-B FUQ for Depression Prediction
Achieving EOC entails two conditions. First, the coverage across different sensitive attribute groups should be equal (we illustrate with groups $s_{1}$ and $s_{2}$, though practical applications may involve more).
$P(y\in C(X^{s})|s=s_{1})=P(y\in C(X^{s})|s=s_{2})\geq 1-\alpha$ (4)
Second, each group should meet the expected coverage. The most straightforward method is to group the data based on sensitive attributes and apply CQR to output confidence intervals separately within each group. While this sounds reasonable, grouping significantly sharpens the available data for each group, thereby diminishing the statistical validity guarantees of UQ.
We leverage a group conditional conformal prediction with equal-mass binning to maximize within-group validity and to facilitate the implementation of fairness interventions [30]. Specifically, it divides all participants in the calibration set into $M$ equal-mass bins, $B_{m}=[l_{m},u_{m})$, $m\in[1,M]$, and $l,u$ denote the lower and upper bounds of the bin respectively, with respect to the label $y$. This bin-dividing method ensures that each bin is given equal importance while maximizing the size of each bin, and it is beneficial to permit the independent optimization of coverage for $M$ bins, thereby ensuring marginal coverage while maintaining the interval widths at their narrowest possible point. Since each bin in group conditional conformal prediction is optimized independently, we still need a fairness-aware optimization for the prediction interval while keeping the interval widths as narrow as possible to be informative. That is
$\min$ $\sum_{m,i\in[1,\mathcal{D}_{cal}]}\frac{|C_{m}(X_{i})|}{|\mathcal{D}_{cal}|}$ (5)
[…]
By leveraging the gradient of coverage with respect to conformal scores, we optimize the confidence intervals across different attributes, resulting in fairer UQ for depression prediction, as shown in Algorithm 1.